# Graph External Attention Enhanced Transformer

### v1

@@ 1 Introduction
Instead of aggregating local neighborhood, graph Transformer captures interaction information between any pair of nodes through a single self-attention layer. Some of the existing works focus on customizing specific attention mechanisms or positional encodings (Dwivedi & Bresson, 2020; Ying et al., 2021; Kreuzer et al., 2021; Hussain et al., 2022; Ma et al., 2023), while others combine message-passing GNNs to design hybrid architectures (Wu et al., 2021; Chen et al., 2022; Rampášek et al., 2022). These methods enable nodes to interact with all other nodes within a graph, facilitating the direct modeling of long-range relations. This may address typical issues such as over-smoothing in GNNs.
While the above-mentioned methods have achieved impressive results, their focus tends to be confined to internal information within the graph, specifically, different nodes or structures in the graph, neglecting potential correlations with other graphs. Indeed, in numerous practical scenarios, such as molecular graph data, strong correlations frequently exist between different graphs (molecules). This is exemplified by the three molecular graphs in Figure 1, all of which possess a benzene ring structure. Under such circumstances, integrating inter-graph correlations into the model can improve the effectiveness of graph representation learning.
@@ 2 Related Work
The Transformer with self-attention, a dominant approach in natural language processing (Vaswani et al., 2017; Devlin et al., 2018), has shown competitiveness in computer vision (Dosovitskiy et al., 2021). Given the remarkable achievements of Transformers and their capability to address crucial challenges of GNNs, graph Transformer has been proposed, attracting increasing attention. Existing works primarily focus on designing tailored attention mechanisms or positional and structural encodings, or combining message-passing GNNs, enabling models to capture complex structures.
A number of works embed topology information into graph nodes by designing tailored attention mechanisms or positional and structural encodings without message-passing GNNs. GT (Dwivedi & Bresson, 2020) introduces the pioneering example of a graph Transformer by integrating Laplacian positional encoding. In the following years, a series of works spring up. SAN (Kreuzer et al., 2021) incorporates both sparse and global attention mechanisms in each layer, utilizing Laplacian positional encodings for the nodes. Graphormer (Ying et al., 2021) attains state-of-the-art performance in graph-level prediction tasks with centrality encoding, spatial encoding and edge encoding. EGT (Hussain et al., 2022) introduces the edge channel into attention mechanisms and adopts an SVD-based positional encoding instead of Laplacian positional encoding. While self-attention is commonly constrained by quadratic complexity, our proposed GEA demonstrates a linear computational complexity with respect to both the number of nodes and edges.
@@ 3.1 Graph External Attention
We first revisit the Transformer, as illustrated in Figure 2(a). Transformer consists of two blocks: a self-attention module and a feed-forward network (FFN). Specifically, self-attention regards the graph as a fully connected graph and computes the attention of each node to every other node. With the input node features $\mathbf{X}$, self-attention linearly projects the input into three matrices: a query matrix ($\mathbf{Q}$), a key matrix ($\mathbf{K}$) and a value matrix ($\mathbf{V}$), where $\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}_{K}$ and $\mathbf{V}=\mathbf{X}\mathbf{W}_{V}$. Then self-attention can be formulated as:
@@
Table 1: Comparison of GEAET to baselines on various tasks, including three tasks in Benchmarking GNNs (graph classification for CIFAR10 and MNIST, node classification for PATTERN), as well as three tasks in LRGB (node classification for PascalVOC-SP and COCO-SP, graph regression for Peptides-struct). Best results are colored: first, second, third.
@@ 3.2 Graph External Attention Enhanced Transformer
Table 2: Comparison of the classical message-passing GNN baseline to its variant augmented by GEANet (without positional encoding for fairness) across five LRGB tasks: (from left to right) node classification, node classification, graph regression, graph classification and link prediction.
@@
Table 3: Comparison of the classical message-passing GNN baseline to its variant augmented by GEANet (without positional encoding for fairness) across five Benchmarking GNNs tasks: (from left to right) node classification, node classification, graph classification, graph classification and graph regression.
@@ 4 Experiments
In summary, our experiments reveal that: (a) our GEAET architecture outperforms existing state-of-the-art methods on various datasets, (b) GEANet can be seamlessly integrated with some basic GNNs, significantly enhancing the performance of GNNs and (c) GEANet shows better interpretability than Transformer and is less dependent on positional encoding.
@@ 4.1 Comparison with SOTAs
As shown in Table 1, for the three tasks from Benchmarking GNNs (Dwivedi et al., 2020), we observe that our GEAET achieves SOTA results on CIFAR10 and MNIST and ranks second on the PATTERN dataset. For the three tasks on LRGB (Dwivedi et al., 2022b), GEAET achieves the best results. It is noteworthy that the GEAET achieves F1 scores of 0.4585 on PascalVOC-SP and 0.3895 on COCO-SP, surpassing other models by a significant gap.
@@ 4.3 Comparison with Self-Attention
024680.2450.250.2550.260.265Number of attention headsTest MAEGCN+GEANetGCN+Transformer
@@
NoneLapPERWPE0.20.20.20.20.250.250.250.250.30.30.30.30.350.350.350.350.3870.3870.3870.3870.2520.2520.2520.2520.2860.2860.2860.2860.2510.2510.2510.2510.2450.2450.2450.2450.2550.2550.2550.255Types of PETest MAEGCN+TransformerGCN+GEANet
Figure 6: Test MAE with different Positional Encodings (PEs).
@@ 4.4 GEAET Mitigates Over-Squashing
The TreeNeighbourMatch dataset (Alon & Yahav, 2021), serves as a synthetic dataset to highlight the issue of over-squashing intuitively. On this dataset, each example represents a binary tree of depth $r$. Figure 7 shows the performance comparison of the standard message-passing GNNs (i.e., GCN, GINE, GAT, GIN and GatedGCN) and our proposed GEAET on the TreeNeighbourMatch dataset. The results indicate that our GEAET generalizes well on the dataset up to $r=7$, effectively alleviating the issue of over-squashing. In contrast, the five GNN methods fail to generalize effectively from $r=4$. This is consistent with the perspective of Alon & Yahav (2021) that GNN suffers from over-squashing due to squashing information of the graph into a fixed-length embedding. In contrast, our method addresses this problem by utilizing graph external attention mechanisms to transmit long-range information without squashing.
@@ 4.5 Ablation Experiments
[…]
Figure 7: Test accuracy across problem radius (tree depth) on the TreeNeighbourMatch dataset.
@@ Appendix A Dataset Descriptions
We evaluate our method on diverse datasets, including five graph benchmark datasets from Benchmarking GNNs, five long-range dependency graph datasets from LRGB and the TreeNeighbourMatch dataset. Below, we provide descriptions of the datasets and present summary statistics in Table 5.
@@
PATTERN and CLUSTER (Dwivedi et al., 2020) are synthetic datasets designed for node classification tasks, both derived from the Stochastic Block Model (SBM; Abbe, 2018) used to model communities. In PATTERN, the objective is to determine whether a node is part of one of the 100 predefined subgraph patterns, while CLUSTER focuses on classifying nodes into six distinct clusters with identical distributions. The unique feature of PATTERN involves recognizing nodes belonging to randomly generated sub-graph patterns, while CLUSTER entails inferring the cluster ID for each node in graphs composed of 6 SBM-generated clusters. We use the splits as is used in (Dwivedi et al., 2020).
@@ Appendix B Hyperparameter Choices and Reproducibility
Table 6: Hyperparameters used for GEAET across datasets: CIFAR10, MNIST, PATTERN, Peptides-Struct, PascalVOC-SP, COCO-SP.
@@ Appendix C Additional Results
Table 7: The results on Transformer and GEANet with different number of attention heads.
@@
Table 8: The results on Transformer and GEANet with different positional encodings.

### v2

@@ 1 Introduction
Instead of aggregating local neighborhood, Graph Transformers (GTs) capture interaction information between any pair of nodes through a single self-attention layer. Some of the existing works focus on customizing specific attention mechanisms or positional encodings (Dwivedi & Bresson, 2020; Ying et al., 2021; Kreuzer et al., 2021; Hussain et al., 2022; Ma et al., 2023), while others combine message-passing GNNs to design hybrid architectures (Wu et al., 2021; Chen et al., 2022; Rampášek et al., 2022). These methods enable nodes to interact with all other nodes within a graph, facilitating the direct modeling of long-range relations. This may address typical issues such as over-smoothing in GNNs.
While the above-mentioned methods have achieved impressive results, they are confined to internal information within the graph, neglecting potential correlations with other graphs. In fact, strong correlations between different graphs generally exist in numerous practical scenarios, such as molecular graph data. Figure 1 shows 3 molecular graphs with a benzene ring structure. Intuitively, exploiting inter-graph correlations can improve the effectiveness of graph representation learning.
@@ 2 Related Work
The Transformer with self-attention, a dominant approach in natural language processing (Vaswani et al., 2017; Devlin et al., 2018), has shown competitiveness in computer vision (Dosovitskiy et al., 2021). Given the remarkable achievements of Transformers and their capability to address crucial challenges of GNNs, GTs have been proposed, attracting increasing attention. Existing works primarily focus on designing tailored attention mechanisms or positional and structural encodings, or combining message-passing GNNs, enabling models to capture complex structures.
A number of works embed topology information into graph nodes by designing tailored attention mechanisms or positional and structural encodings without message-passing GNNs. GT (Dwivedi & Bresson, 2020) is the pioneering work of GTs by integrating Laplacian positional encoding. In the following years, a series of works spring up. SAN (Kreuzer et al., 2021) incorporates both sparse and global attention mechanisms in each layer, utilizing Laplacian positional encodings for the nodes. Graphormer (Ying et al., 2021) attains state-of-the-art performance in graph-level prediction tasks with centrality encoding, spatial encoding and edge encoding. EGT (Hussain et al., 2022) introduces the edge channel into attention mechanisms and adopts an SVD-based positional encoding instead of Laplacian positional encoding. While self-attention is commonly constrained by quadratic complexity, our proposed GEA demonstrates a linear computational complexity with respect to both the number of nodes and edges.
@@ 3.1 Graph External Attention
We first revisit the Transformer, as illustrated in Figure 2(a). Transformer consists of two blocks: a self-attention module and a feed-forward network (FFN). Specifically, self-attention regards the graph as a fully connected graph and computes the attention of each node to every other node. With the input node features $\mathbf{X}$, self-attention linearly projects the input into 3 matrices: a query matrix ($\mathbf{Q}$), a key matrix ($\mathbf{K}$) and a value matrix ($\mathbf{V}$), where $\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}_{K}$ and $\mathbf{V}=\mathbf{X}\mathbf{W}_{V}$. Then self-attention can be formulated as:
@@
Table 1: Comparison of GEAET with baselines on 6 datasets. Best results are colored: first, second, third.
@@ 3.2 Graph External Attention Enhanced Transformer
Table 2: Comparison of the classic message-passing GNN baselines with their variants augmented by GEANet on 5 benchmarks.
@@
Table 3: Comparison of the classic message-passing GNN baselines with their variants augmented by GEANet on 5 benchmarks.
@@ 4 Experiments
In summary, our experiments reveal that (a) GEAET architecture outperforms existing state-of-the-art methods on various datasets, (b) GEANet can be seamlessly integrated with some basic GNNs, significantly enhancing the performance, (c) GEANet shows better interpretability than Transformer and (d) GEANet is less dependent on positional encoding.
@@ 4.1 Comparison with SOTAs
As shown in Table 1, for the 3 tasks from Benchmarking GNNs (Dwivedi et al., 2020), we observe that our GEAET achieves SOTA results on CIFAR10 and MNIST and ranks second on the PATTERN dataset. For the 3 tasks on LRGB (Dwivedi et al., 2022b), GEAET achieves the best results. It is noteworthy that the GEAET achieves F1 scores of 0.4585 on PascalVOC-SP and 0.3895 on COCO-SP, surpassing other models by a significant gap.
@@ 4.3 Comparison with Self-Attention
024680.2450.250.2550.260.265Number of attention headsMAEGCN+GEANetGCN+Transformer
@@
NoneLapPERWPE0.20.20.20.20.250.250.250.250.30.30.30.30.350.350.350.350.3870.3870.3870.3870.2520.2520.2520.2520.2860.2860.2860.2860.2510.2510.2510.2510.2450.2450.2450.2450.2550.2550.2550.255Types of positional encodingMAEGCN+TransformerGCN+GEANet
Figure 6: Test MAE with different positional encodings.
@@ 4.4 GEAET Mitigates Over-Squashing
The TreeNeighbourMatch dataset (Alon & Yahav, 2021) is used to provide an intuition of over-squashing. On this dataset, each example is represented with a binary tree of depth $r$. Figure 7 shows the performance comparison of the standard message-passing GNNs with our proposed GEAET on the TreeNeighbourMatch dataset. The results indicate that our GEAET generalizes well on the dataset up to $r=7$, effectively alleviating the issue of over-squashing. In contrast, the 5 GNN methods fail to generalize effectively from $r=4$. This is consistent with the perspective of Alon & Yahav (2021) that GNN suffers from over-squashing due to a fixed-length embedding of the graph. Our method addresses this problem by utilizing graph external attention mechanisms to transmit long-range information.
[…]
Figure 7: Test accuracy with different problem radius $r$ on the TreeNeighbourMatch dataset.
@@ Appendix A Dataset Descriptions
We evaluate our method on diverse datasets, including 5 graph benchmark datasets from Benchmarking GNNs, 5 long-range dependency graph datasets from LRGB and the TreeNeighbourMatch dataset. Below, we provide descriptions of the datasets and present summary statistics in Table 5.
@@
PATTERN and CLUSTER (Dwivedi et al., 2020) are synthetic datasets derived from the Stochastic Block Model (SBM; Abbe, 2018) on node classification tasks. PATTERN is to determine whether a node belongs to one of the 100 predefined subgraph patterns, while CLUSTER is to classify nodes into 6 distinct clusters with identical distributions. The unique feature of PATTERN involves recognizing nodes belonging to randomly generated sub-graph patterns, while CLUSTER entails inferring the cluster ID for each node in graphs composed of 6 SBM-generated clusters. We use the splits as is used in (Dwivedi et al., 2020).
@@ Appendix B Hyperparameter Choices and Reproducibility
Table 6: Hyperparameters used for GEAET on 6 datasets.
@@ Appendix C Additional Results
Table 7: The results of Transformer and GEANet with different number of attention heads.
@@
Table 8: The results of Transformer and GEANet with different positional encodings.